Hello, I need help with the following word problem:

Question:

"You are a detective trying to solve a murder. A body was found by a janitor with an initial temperature of 85 degrees Fahrenheit at 7:00 AM. One hour later, when you arrive at 8:00 AM, the body of the victim has dropped to 80 degrees Fahrenheit. The temperature of the office was at a constant 70 degrees Fahrenheit. With your knowledge of the human anatomy, you know that the body temperature at the time of death had to be 98.6 degrees Fahrenheit. With all the facts and evidence before you, what was the time of death for the victim?

Hint: Using your detective senses and a little bit of math, you know to use the following formula to help you find the first clue to the time of death:

dT d = k0T-TA."

Thank your time and help!
Answer #1

Ath the temperature of body at the times then dI = (T- Tom) = (-70) (Here Tom=707) * log(T-70) = k++ c at t=0, T=98.6 → C= lo